Break-Even Sales and A chart used to assist management in understanding the relationships among costs, expenses, sales, and operating profit or loss.Cost-Volume-Profit Chart

Last year, Gelbin Inc. had sales of $209,000, based on a unit selling price of $110. The Costs that vary in total dollar amount as the level of activity changes.variable cost per unit was $80, and Costs that tend to remain the same in amount, regardless of variations in the level of activity.fixed costs were $39,600. The maximum sales within Gelbin's The range of activity over which changes in cost are of interest to management.relevant range are 2,400 units. Gelbin is considering a proposal to spend an additional $12,000 on billboard advertising during the current year in an attempt to increase sales and utilize unused capacity.

Required:

1. Construct a cost-volume-profit chart on your own paper, indicating the break-even sales for last year. In your computations, do not round the contribution margin percentage.

Break-Even sales (dollars)

Break-even sales (units)

2. Using the cost-volume-profit chart prepared in part (1), determine (a) the income from operations for last year and (b) the maximum income from operations that could have been realized during the year. In your computations, do not round the contribution margin percentage.

Income from operationf

Maximum income from operations

3. Construct a cost-volume-profit chart (on your own paper) indicating the break-even sales for the current year, assuming that a noncancelable contract is signed for the additional billboard advertising. No changes are expected in the unit selling price or other costs. In your computations, do not round the contribution margin percentage.

Dollars

Units

4. Using the cost-volume-profit chart prepared in part (3), determine (a) the income from operations if sales total 1,900 units and (b) the maximum income from operations that could be realized during the year. In your computations, do not round the contribution margin percentage.

Income from operations at 1,900 units

Maximum income from operations
Answer #1

Question 1

A cost volume profit chart can be drawn by plotting the sales, fixed costs and variable costs at different sales volumes.

From the the above info we can infer the unit price to be $110. Since we are drawing the CVP chart for the previous year, we ignore the billboard advertising expense budgeted for this year. We use the following data to plot the chart

Units Sold Variable Cost (80XUnits) Fixed Costs Sales (110 X Units)
500                    40,000                    39,600                    55,000
600                    48,000                    39,600                    66,000
700                    56,000                    39,600                    77,000
800                    64,000                    39,600                    88,000
900                    72,000                    39,600                    99,000
1000                    80,000                    39,600                 1,10,000
1100                    88,000                    39,600                 1,21,000
1200                    96,000                    39,600                 1,32,000
1300                 1,04,000                    39,600                 1,43,000
1400                 1,12,000                    39,600                 1,54,000
1500                 1,20,000                    39,600                 1,65,000
1600                 1,28,000                    39,600                 1,76,000
1700                 1,36,000                    39,600                 1,87,000
1800                 1,44,000                    39,600                 1,98,000
1900                 1,52,000                    39,600                 2,09,000
2000                 1,60,000                    39,600                 2,20,000
2100                 1,68,000                    39,600                 2,31,000

Using the above data, we can plot the following graph:

Using the above chart we can derive the required figures:

contribution margin is Sales Price - Variable Cost = 110 - 800 = 30

contribution margin % = 30/110 = 27.2727%

Break Even Sales is computed as Fixed Costs/Contribution = 39600/30 = 1320 units

This in dollars can be computed as 1320*110 = $145,200

Question 2

Using the cost volume profit chart we can obtain the income from operations as

1900*30 - 39600 = $17,400

The maximum income from operations can be computed as

2400*30 -39600 = 32400

Question 3

In case of the current year, we add the billboard cost of 12000 to the fixed costs to compute the CVP Chart. Nothing else changes:

Units Sold Variable Cost (80XUnits) Fixed Costs Sales (110 X Units) Total Cost (Fixed + Variable) Profit
500                    40,000                    51,600                    55,000                    91,600     -36,600
600                    48,000                    51,600                    66,000                    99,600     -33,600
700                    56,000                    51,600                    77,000                 1,07,600     -30,600
800                    64,000                    51,600                    88,000                 1,15,600     -27,600
900                    72,000                    51,600                    99,000                 1,23,600     -24,600
1000                    80,000                    51,600                 1,10,000                 1,31,600     -21,600
1100                    88,000                    51,600                 1,21,000                 1,39,600     -18,600
1200                    96,000                    51,600                 1,32,000                 1,47,600     -15,600
1300                 1,04,000                    51,600                 1,43,000                 1,55,600     -12,600
1400                 1,12,000                    51,600                 1,54,000                 1,63,600       -9,600
1500                 1,20,000                    51,600                 1,65,000                 1,71,600       -6,600
1600                 1,28,000                    51,600                 1,76,000                 1,79,600       -3,600
1700                 1,36,000                    51,600                 1,87,000                 1,87,600          -600
1800                 1,44,000                    51,600                 1,98,000                 1,95,600        2,400
1900                 1,52,000                    51,600                 2,09,000                 2,03,600        5,400
2000                 1,60,000                    51,600                 2,20,000                 2,11,600        8,400
2100                 1,68,000                    51,600                 2,31,000                 2,19,600      11,400
2400                 1,92,000                    51,600                 2,64,000                 2,43,600      20,400

The new CVP looks like this:

Since there is no change in variable costs, the contribution Margin does not change. However the new break even point is:

51600/30 = 1720 Units.

BEP in Dollars is 1720 X 110 = $189,200.

Question 4

The income of the company at 1900 sales is

1900 X 30 - 51600 = $5,400

The Maximum income at 2400 Units is

2400 X 30 - 51600 = 20,400